OpenStudy (anonymous):

My sister wants to know what 8*4=So i told her she can use my openstudy.com account.Plz tell her what 8 times 4 equals?

TheSmartOne (thesmartone):

Calculator,

OpenStudy (abcdefghijklmq):

8*4=32

OpenStudy (anonymous):

See sis it equal 32.

TheSmartOne (thesmartone):

\(\sf 8 \times 4 = 8+ 8 +8+8\)

OpenStudy (anonymous):

AWWWWWW so cute lol who helpful

OpenStudy (anonymous):

My sis says (Thanks)

TheSmartOne (thesmartone):

Or it equals to: \(\sf 8 \times 4 = 4 + 4 + 4 + 4 + 4 + 4+4 +4\)

OpenStudy (anonymous):

Thanks for helping my sis out guys

OpenStudy (abcdefghijklmq):

no problem

TheSmartOne (thesmartone):

https://www.google.com/?gws_rd=ssl#q=calculator